OpenStudy (gg):

solve ordinary differential equation (1-x^2)y'-2xy^2=xy

OpenStudy (jamesj):

This equation is separable; see how to do that and then you're in the home stretch.

OpenStudy (anonymous):

James could you kindly please go through the workings, i am studying this at the moment as well.

OpenStudy (jamesj):

We want to write y' = F(x,y) = g(x)h(y)

OpenStudy (gg):

I got \[dy/(2y ^{2}-y)=xdx/(1-x ^{2})\]

OpenStudy (jamesj):

Because if we can, then \( dy/h(y) = g(x) dx \) and we can integrate.

OpenStudy (jamesj):

Yes, almost right. Check your signs carefully.

OpenStudy (gg):

I found a mistake :)

OpenStudy (jamesj):

\[ (1-x^2) y' = x (y + 2y^2) \]

OpenStudy (gg):

and then we express y from the result and that's it?

OpenStudy (jamesj):

yes.

OpenStudy (gg):

thanks :) can I give some problems more? can I attach a file?

OpenStudy (jamesj):

For example, the integral of the x side is \( -(1/2) \ln(1-x^2) + C \) No .... I'm off to buy my new iPhone right now.

OpenStudy (jamesj):

Sorry ;-)

OpenStudy (gg):

u can do it later? :)

OpenStudy (jamesj):

I'll probably be around later, yes.

OpenStudy (gg):

ok :) thanks :) good luck in shopping ;)